OpenStudy (anonymous):

write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1,1)

OpenStudy (lgbasallote):

\[y - y_1 = m(x-x_1)\] \[y - 1 = 6(x-1)\] \[y - 1 = 6x - 6\] put it now in the slope-intercept form